### v1

# LISTN: Lexicon induction with socio-temporal nuance

Abstract
Research on extremist online communities frequently utilizes linguistic analysis to explore group dynamics and behaviour. Existing studies often rely on outdated lexicons that do not capture the evolving nature of in-group language, nor the social structure of the community. This paper proposes a novel method for inducing in-group lexicons which incorporates its socio-temporal context. Using dynamic word and user embeddings trained on conversations from online anti-women communities, our approach outperforms prior methods for lexicon induction. We provide a new lexicon of manosphere terms, validated by human experts, which quantifies the relevance of each term to a specific sub-community. We present novel insights on in-group language which illustrate the utility of this approach.
1 Introduction
Language cues are often used to investigate extremist movements (Baele et al., 2023; Vergani and Bliuc, 2015) or to analyse the behaviour of group members (Fernandez et al., 2018). One reason for this is that groups have a tendency to develop distinctive vernaculars, also known as in-group language. According to Drake (1980), in-group language plays two important social roles within groups: first, it can obscure the discussions of the group from out-group onlookers or moderators, which is of particular importance to deviant and extremist online communities (Chancellor et al., 2016); and second, it signals cohesion and solidarity with a group by illustrating an awareness and acceptance of their norms. Since this type of informal language evolves rapidly (Stewart and Eisenstein, 2018), the use of recent lexical innovations of the group is a strong signal of group belonging (Danescu-Niculescu-Mizil et al., 2013; Zhang et al., 2017).
Studies that use hand-compiled lexicons are dependent on the skill of the lexicon builders (Havaldar et al., 2024). In the case of in-group language, there are subtle signals that are challenging to identify as an out-group member (Mendelsohn et al., 2023), especially given the scale of modern text corpora. Since manual lexicon building is expensive, there is a tendency not to update lexicons frequently. For example, Rowe and Saif (2016) include a 7 year old lexicon in their study on online pro-ISIS language. These factors impact the value of the lexicon with resulting negative effect on the quality of downstream studies.
Despite the highly dynamic and social nature of groups and their language, existing methods for computationally inducing in-group language tend to focus only on linguistic information. For example, Lucy and Bamman (2021) use statistical properties of word occurrences and contextualised word embeddings. Similarly, Farrell et al. (2020) and Yoder et al. (2023) use approaches based on topic models and static word embeddings.
In this work, we propose methods for lexicon induction with socio-temporal nuance (LISTN). Our approach leverages dynamic word and user embeddings, developed in a recent work by De Kock (2024), in the context of online anti-women communities collectively known as the manosphere. Our experiments indicate that LISTN-based methods outperform a range of baselines based on prior work, achieving an average precision score of 0.77 on a newly crafted test set for lexicon induction. We use existing lexicons of manosphere language and human expert validation in our evaluation, and release a lexicon of 455 new manosphere terms with scores representing their relevance to specific manosphere subcommunities.
Using this lexicon, we further find that word embeddings of in-group terms tend to be more static (i.e. showing less variation across timesteps) compared to words of similar frequencies in the general vocabulary. We further investigate the interaction of in-group language and violent language in the manosphere, finding that there is a positive correlation ($\rho=0.53$) but that the strength of the relationship varies for different subcommunities. Finally, we look at the overlap in vernacular between the different groups, finding that Incels are the most distinctive in their in-group language. Our findings provide quantitative support for prior claims in sociological studies of relationships between different sub-groups of the manosphere.
2 Socio-temporal models of language
As mentioned in Section 1, we leverage word and user embeddings from prior work. The architecture of Yao et al. (2018) deconstructs user-content representations over different timesteps, which yields dynamic user embeddings and word embeddings. A recent work by De Kock (2024) extended this model, proposing a joint model for community structure and language, which accounts for temporal evolution in both factors. They show that the resulting embeddings capture finegrained changes in the relevance of a term to a community, and to sub-groups within the community.
To induce lexicons with such a joint model, we use pretrained embeddings for both of the abovementioned architectures, as produced by De Kock (2024), in this work. The models are pretrained on data from the manosphere, which is known to have highly distinctive and dynamic in-group language (Bogetić, 2023; Farrell et al., 2020) and sub-group structure (Ribeiro et al., 2021), making it well-suited for our focus on socio-temporal context. Below we review the architectures and the training data used. Full details are provided in the respective papers.
Content-only
A user-content matrix $C_{t}$ is constructed, using pointwise mutual information (PMI) to capture how frequently a user uses a given word compared to a background corpus during timestep $t$. Matrix factorisation is used to obtain user embeddings $U_{t}$ and word embeddings $W_{t}$ such that $C_{T}\approx U_{t}\cdot W_{t}^{T}$, for all $t$. To ensure alignment of embeddings across timesteps, the learning objective includes a regularisation term that penalises large inter-timestep changes.
Content and adjacency
To capture interpersonal dynamics and social structure, a user-user adjacency matrix $A_{t}$ is constructed to capture how frequently two users interact in the same thread at time $t$. This matrix is jointly decomposed with $C_{t}$ to yield $U_{T}$, $V_{T}$ and $W_{T}$ such that $C_{T}\approx U_{t}\cdot W_{t}^{T}$ and $A_{T}\approx U_{t}\cdot V_{t}^{T}$. A similar smoothing mechanism is used to align embeddings over time. De Kock (2024) show that the joint factorisation outperforms the content-only approach in their evaluation.
Data
The models are trained on over 4 million utterances from 50 manosphere subreddits over 9 months, using the Reddit portion of the corpus of Ribeiro et al. (2021). The subreddits in the dataset are mapped to five different manosphere sub-communities, referred to as categories in this paper: Incels, Men’s Rights Activists (MRA), Men Going Their Own Way MGTOW, Pick-up Artists (PuA), and The Red Pill (TrP). The dataset further includes data from two anti-manosphere communities (referred to as Criticism) and three Mental Health related communities, which we use as a control group in this study.
Embeddings are produced at monthly timesteps, representing April-December 2018. The data is filtered to include only users who are active for 3 or more timesteps and words that are used by more than 20 users. The PMI in the $C$ matrix is calculated against a background corpus of 40 million Reddit posts by Dziri et al. (2019) is used. For tokenisation, the NLTK word tokeniser is used. Words are lowercased and non-alphabetic characters are removed.
3 LISTN
Using the models described in Section 2, we evaluate a number of approaches for inducing lexicons. We refer to these methods as LISTN: lexicon-induction with socio-temporal nuance, and distinguish between LISTN-CA (jointly training on content and adjacency matrices) and LISTN-C (using only the content matrix).
Recall that the user-content matrix $C$ is constructed using PMI to compare how often a user uses a term compared to a large background corpus. To use the resulting word embeddings $w$ and user embeddings $u$ for lexicon induction, we evaluate several methods based on the low-rank reconstruction $u_{t}\cdot w_{t}^{T}$, referred to as the relevance of a word to a user by De Kock (2024). The rationale is that we hope to find words which are used more often than expected (as indicated by PMI), using embeddings to provide a more nuanced view of a word or user compared to word frequencies. Farrell et al. (2020) find that the usage of in-group terms varies between the different manosphere subcommunities.
Therefore, we explore several methods that allow for sub-group specialisation.
Community centroid:
We calculate a vector of word relevance scores $y_{t}=\overline{u}_{t}\cdot w_{t}^{T}$, where $\overline{u}_{t}$ represents the centroid of all users in the training set at time $t$.
Category:
We calculate the centroid of each of the manosphere categories (as described in Section 3) $\overline{c}_{t}$ as the mean of all users who contributed to that category at timestep $t$. For every word, we then find its maximum relevance score across all categories: $y_{t}=\max_{c_{t}}(\overline{c}_{t}\cdot w_{t}^{T})$. We use maximum aggregation (as opposed to mean or median) because a word only needs to be highly relevant to one sub-group to be considered as in-group language.
Subreddit:
Similar to the category-based method, we find the maximum relevance score at the subreddit level using its centroid $\overline{s}_{t}$: $y_{t}=\max_{s_{t}}(\overline{s}_{t}\cdot w_{t}^{T})$.
Cluster:
We use K-means clustering of the user embeddings to find subcommunities. The motivation is that there may be groups that span different categories or subreddits, or sub-groups within a given category. For each cluster $r$ we find $y_{t}=\max_{r_{t}}(\overline{r}_{t}\cdot w_{t}^{T})$. We do this for $K\in\{5,20,100\}$. Since the clustering dependends on random initialisation, we report the mean over 5 runs.
Bootstrap:
Departing from the reconstruction-based approaches, we use a list of confirmed terms from manosphere lexicons (more details in Section 4.3) and find their nearest neighbours based on cosine similarity. The score for each neighbouring word is given by its maximum similarity to a lexicon word.
Bias:
The factorisation models described in Section 2 include an inferred bias term for each user and word to capture population-level trends. In the case of the word bias, it captures the tendency for a word to have a high PMI across all users as compared to a background corpus. We use this directly as the word score.
These approaches produce a score for each term at every timestep. To aggregate, we use the maximum score of a term across all timesteps. This has the benefit of potentially capturing words that move into and out of popularity over the training window.
4 Experimental setup
4.1 Task definition
The goal of this work is to induce lexicons of manosphere language. For the purpose of evaluation, we focus on single-token lexical innovations. Our rationale for this task definition is provided below.
A challenge in evaluating lexicons is that they are not a precise science and often rely on subjective expert judgment, as discussed in Section 1. For example, the lexica in Table 3 contain words such as hurt and dumb, which represent borderline cases in deciding whether they are sufficiently relevant for inclusion. A further challenge is that dogwhistles (words that take on a special meaning within an in-group that is different from its meaning to the general population) are difficult for researchers to identify (Lucy and Bamman, 2021; Mendelsohn et al., 2023).
To address these challenges, we focus on lexical innovations in our evaluation. Lexical innovations are a form of in-group language which involve non-standard or extragrammatical words (Slotta, 2016); for example, “foid” is a common derogatory term for a woman in the Incel community. Lexical innovations provide a clearer decision boundary as they tend to be more easily attributed to a specific community. Notably, this is a constraint on the evaluation and not a limitation of the method; all methods used in this work can produce scores for standard and non-standard language alike.
A set of standard words are required in order to identify non-standard words, or lexical innovations. For this purpose, we use the vocabulary of GloVe-400k (Pennington et al., 2014), which is based on news and Wikipedia articles. We make an exception to include in-vocabulary words from existing manosphere lexicons in our evaluation, irrespective of whether they are lexical innovations. Since prior work has deemed these terms to be of particular importance to the manosphere, we consider them to be relevant and useful for evaluation.
Finally, we consider only single-token words to limit the vocabulary size. Future work may investigate N-gram models, since multi-word in-group expressions are also observed in the data. However, the common multi-word expressions are often made into acronyms; for example, AWALT is an acronym of “all women are like that”.
4.2 Baselines
We use baselines based on two prior lexicon induction approaches:
Embedding expansion
We train a word2vec model (Mikolov et al., 2013) on the same subset of posts used to train the LISTN models. We then use a bootstrap method to find the nearest neighbours of a lexicon of known manosphere terms using cosine similarity. Similar approaches are followed by Yoder et al. (2023), Farrell et al. (2020) and Havaldar et al. (2024) for lexicon expansion.
Statistical measures
Similar to Lucy and Bamman (2021) and Zhang et al. (2017), we evaluate variants of PMI, including normalised PMI (NPMI). Given a word $w$ and a particular setting (e.g. subreddit or community) $s$:
$\text{PMI}(w,s)=\text{log}\frac{P(w|s)}{P(w)},$ (1)
$\text{NPMI}(w,s)=\frac{\text{PMI}(w,s)}{-\text{log}P(w,s)}.$ (2)
Probabilities are calculated based on word frequencies. To determine $P(w)$ and $P(w,s)$, we use the same background corpus used to construct the $C$ matrix. The (N)PMI of a term is then calculated based on its probability in our training set compared to the background corpus. To account for potential sub-group specialisation, we also evaluate (N)PMI on a subreddit and category level. We use the maximum PMI for a given word across the set of subcommunities to produce its score. Finally, we evaluate the (N)PMI of the full community using monthly slices of data and maximum aggregation.
4.3 Existing lexicons
We use existing lexicons of manosphere language as basis for evaluating our methods. The lexicons and the number of terms, filtered to include only terms that are in our corpus, are shown in Table 1.
Source # words
IncelWiki [^1] https://incels.wiki/w/Incel_Glossary 189
Farrell et al. (2019) 194
Baele et al. (2023) 122
Lucy and Bamman (2021) 52
Hatebase [^2] https://hatebase.org/search_results/gender_id%3D6%7Clanguage_id%3Deng 29
Total (deduplicated) 483
Table 1: Pre-existing lexicons used in the evaluation, and the number of terms from each lexicon that occur in the training data.
For the Farrell et al. (2019) dataset, we follow Ribeiro et al. (2021) in excluding the lexicon categories that are not specific to the manosphere (e.g. racism). The Hatebase and IncelWiki lexica are crawled by us and will be released to the community. Combined, this provides a set of 483 terms.
4.4 Evaluation
To evaluate our methods, we first perform a preliminary scoring step that is used to construct a test set. This test set is used as gold standard in our main evaluation.
Preliminary scoring
In this step, we evaluate each of the approaches in Sections 3 & 4.2 against the lexicons in Section 4.3. Terms in the lexicons are assigned a True label and all others are considered False. An issue with this approach, however, is that the lexicons suffer from low recall (a known issue; see Lucy and Bamman, 2021), meaning that true in-group terms identified by the methods are often evaluated as false positives. For this reason, we use this scoring only to identify terms to include in our test set (described below). The average precision and area under the receiver-operator characteristic curve of each method are provided in Appendix A.
Test set construction
We retrieve the top-scoring 1000 terms produced by the best baseline and best LISTN approaches from the preliminary scoring step. These terms are manually evaluated by an social psychology PhD student who is a specialist on the manosphere, as well as an author of this paper. Annotators are tasked with assigning a binary label to each word, based on whether or not it strongly signals membership of the manosphere. They are also provided with three example usages of each word and the name of the subreddit where it was observed. Following the expert annotator’s advice, we also annotate “cultural exports” – terms that originated within the manosphere but have spread to be used more widely. In our evaluation, we treat these terms as positives. We further include the terms from the existing lexicons in Table 1 as positive samples.
The resulting test set consists of 1803 words with binary labels, with an inter-annotator agreement score of 0.726 using Cohen’s Kappa, indicating substantial agreement. In cases of disagreement, we revert to the label from the expert annotator. The test set is reasonably balanced, consisting of 944 positive and 859 negative samples. This dataset will be released to the public.
Main evaluation
This dataset is used for evaluation in all experiments in this paper. We report the average precision (AP) and area under the receiver-operator characteristic curve (AUROC) for each method. The AP provides a measure of the model’s precision over all possible recall values; that is, how well the model avoids false positives. AUROC is a measure of its ability to distinguish between positive and negative cases. Davis and Goadrich (2006) show that a model dominates in AUROC space if and only if it dominates in AP space, whereas the converse is not true; as such, we prioritise the AP. We use the randomised permutation test with Monte Carlo approximation ($R=9999$ and $\alpha=0.05$) for significance testing.
5 Results
AP AUROC
Random 0.52 0.5
Baselines
Word2Vec bootstrap 0.5563 0.5427
PPMI 0.6333 0.6709
PPMI-subreddit 0.5910 0.5916
PPMI-category 0.6517 0.6470
PPMI-month 0.6165 0.6071
NPMI 0.6709 0.7034
NPMI-subredit 0.6170 0.6187
NPMI-category 0.6790 0.6647
NPMI-month 0.6449 0.6386
LISTN-CA
Community centroid 0.6228 0.5131
Category centroid 0.7231 0.6713
Subreddit centroid 0.5519 0.5043
Cluster-5 0.7620 0.7403
Cluster-20 0.7069 0.6954
Cluster-100 0.6950 0.6566
Bootstrap 0.5349 0.5206
Bias 0.6190 0.5560
LISTN-C
Community centroid 0.6297 0.5255
Category centroid 0.7272 0.6809
Subreddit centroid 0.5891 0.5461
Cluster-5 0.7679 0.7363
Cluster-20 0.7554 0.7289
Cluster-100 0.7040 0.6864
Bootstrap 0.5276 0.5077
Bias 0.6016 0.5383
Table 2: Comparison of different lexicon induction approaches.
Results are shown in Table 2. The best LISTN-CA and LISTN-C methods both outperform the best baseline method (statistically significant, $P<<0.05$). Using LISTN-C produces a higher AP score compared to LISTN-CA for 6 out of the 8 induction methods, including the top-performing method (not statistically significant; $P=0.723$). This is a surprising result, as it indicates that the induction does not benefit from the added author adjacency information (i.e. shared thread interaction) in LISTN-CA. A benefit of this outcome is that the LISTN-C model has fewer parameters and is more computationally efficient to train. The LISTN-C model does still contain social information, in that the source matrix is a user-content representation, meaning that users are represented as being similar if they use the same words. By comparison, the standard word2vec matrix factorisation formulation only considers linguistic context – words are similar if they co-occur with the same context words (Goldberg and Levy, 2014).
LISTN
For both LISTN-CA and LISTN-C, the 5-cluster approach performs the best. Notably, it performs better than category-level induction, which computes centroids of a similar level of granularity, indicating that the embedding clustering is identifying useful sub-groups which are not evident from the subreddit structure. Using more clusters ($K=20$ and $K=100$) does not improve performance. A possible explanation might be that smaller clusters lead to a less precise or less informative centroid which overemphasises irrelevant words. Subreddit-level induction (a more granular version of the category-level method) performs the worst of the centroid-based approaches for both models and metrics. However, averaging over the full community performs the second worst, indicating that it is necessary to take into account some level of sub-group linguistic specialisation.
Baselines
All baseline approaches outperform the random baselines. The worst-performing baseline is the word2vec-based method. For both LISTN models, the embedding expansion approaches (“bootstrap”) perform the worst as well. A possible reason for this might be that the seed lexicons are not varied enough, such that they are missing some area of the embedding space. The NPMI baseline performs the best on AUROC, whereas the category-level NPMI performs the best for the AP. Generally, the two measures tend to co-vary. In all cases, the NPMI outperforms the PPMI approaches, which corresponds with the results of Lucy and Bamman (2021). Interestingly, using month-level NPMI scoring does not improve over the time-aggregated NPMI, indicating that naive models for accounting for time are not sufficient.
6 Analysis
Beyond the improved performance of lexicon induction methods (as illustrated in Section 5), an advantage of this approach is that it produces temporal embeddings for words and for users. In this section, we use these embeddings to investigate changes in words over time and sub-group specialisation of in-group language in the manosphere.
6.1 Change in word representations
Figure 1: The relationship between word frequency and embedding stability .
Recall that the training objective includes a regularisation term that penalises changes between successive word representations, meaning that changes in embeddings are only permitted if they reduce the reconstruction error sufficiently. In other words, word embeddings change as much as is needed to represent the data, but no more. To determine how much a word changes between two successive timesteps, we use the cosine distance between the embeddings. Its sum over all timesteps is calculated as a measure for how much the word representation varies over the training window.
Using this metric, we find that the words with the lowest variation are stopwords, e.g. from, here, there. This makes sense as the embeddings are a function of how a word is used throughout a community, and these words are not expected to be impacted by the social evolution of the group. At the opposite end of the spectrum, we observe less common words, including betta, lackeys, wooo, fragment, thirstiest, and booklet. Noting that there seems to be a relation between word frequency and embedding variation, we calculate their Spearman correlation and find a strongly negative correlation, $\rho=-0.77$ ($P<<0.05$).
Figure 1 shows the relation between the log frequency of a word and its cumulative embedding variation, for all words in the training data (grey dots). The strong negative correlation only appears to hold for frequencies under approx. 10 000; beyond this point, changes become close to zero. We also observe that the variance in the amount of change covaries with the log frequency, such that a large range of values is observed for low-frequency words. We have two hypotheses for why this might be. Firstly, it may simply be a computational issue, whereby words that are less common do not have high-quality representations due to insufficient training data, and therefore their embeddings are less stable. The second option is that words that are less common truly have more pliability in their colloquial usage, and are used in more subtle, differentiated ways. We expect the reality is a combination of the two.
To investigate this idea, we also plot the confirmed in-group language along the same axes (colourful dots) in Figure 1. Dotted lines indicate straight-line regressions to the whole population and to the in-group language sub-population; we observe that the regression to the in-group population is less steep. This seems to suggest that in-group language terms are more stable and less pliable in how they are used within this community – providing some support for the idea that the variation is only partly explained by frequency. The different colours represent the predictions made by the best models in Section 5, showing the top 1000 words predicted by each approach. The two methods appear to favour different areas in the frequency distribution, with the NPMI-based model being biased to the lower end of the spectrum, and the LISTN-based model to the mid and higher side. Future work may investigate combining these methods for lexicon induction, as they appear to have different strengths.
Further investigating the low-frequency, low-variation terms, we find that it includes several mentions of medications (e.g. lamictal, pristiq, seroquel). These are concepts with very specific, technical meanings that are not likely to vary due to socio-temporal changes in the community. We also note the term thotaudit, which refers to the hashtag of a harassment campaign targeting sex workers [^3] https://www.vice.com/en/article/thotaudit-databases-of-sex-workers-and-reporting-them-to-paypal/. On the other end of the spectrum, the terms submission and bimbo display a large amount of variation relative to their frequency. A possible reason for this might be that these terms occupy different niches for the subcommunities of the manosphere, or that they underwent a socio-semantic change during the training period. While more exploration of this embedding space is warranted, it is evident that it captures some interesting and useful qualities of the data.
6.2 Correlation with violent language
Figure 2: Relevance of violence lexicon and in-group lexicon to different subreddits.
In this section, we explore how violence interacts with in-group language, using the lexicon of confirmed manosphere language from Section 5. Chancellor et al. (2016) found that lexical variation is associated with more extreme content in pro-anorexia communities, but this has not been explored in the context of the manosphere.
Similar to De Kock (2024), we calculate centroids for each of the 50 subreddits in the dataset as the mean of all users who participated in that subreddit at a given timestep. Then, the relevance score of the lexicon is given by the dot product of the lexicon centroid with each of the sub-community centroids. For each subreddit, a score is calculated at each timestep, since the community and word embeddings are both dynamic. Figure 2 shows the relevance scores for the manosphere lexicon (top) and the violence lexicon (bottom). Community sizes are reflected by the relative sizes of the different dots.
Based on this characterisation, Incel subreddits display the highest relevance scores for the manosphere lexicon, with the exception of the r/foreveralonedating and r/foreveralone subreddits, who also have the lowest violence scores of the incel groups. Although the taxonomy of Ribeiro et al. (2021) placed this subreddit in the Incels category, Bachaud (2024) states r/foreveralone actively distances itself from Incels, which is supported by our analysis.
The Pick-up Artists, who have relatively low violence scores, also have relatively low in-group language scores. On the other hand, the MGTOW and MRA groups have higher violence scores but relatively low in-group language scores, showing that these metrics are not always correlated. Calculating the Spearman correlation between the violence and in-group language scores of all users in the dataset, we find a moderate positive correlation of $\rho=0.53$, $(P<<0.05)$. Breaking this down per manosphere category, we observe the highest correlation ($\rho=0.65$) for the Incels, and the lowest for MRA ($\rho=0.45$).
The relevance of the manosphere lexicon is the lowest for the mental health subreddits, which is consistent with our expectations for the control group. The r/suicidewatch group has an above average violence score, whereas the r/depression subreddit is on par with the community average.
6.3 Sub-group linguistic specialisation
The analysis in Section 6.2 indicates that there is a complex relationship between the utilisation of in-group language (as defined by our lexicon) and violent language, which differs at the sub-group level. This, again, points to the existence of sub-group linguistic specialisation among manosphere groups, as mentioned in Section 1. In this section, we further investigate this idea. The hypothesis is as follows: if manosphere sub-groups share a similar vernacular but express it to varying degrees, we would expect there to be a (nonlinear) correlation between their utilisation of different words in the lexicon. On the other hand, if the groups favour disjoint parts of the lexicon, there would be no such correlation.
To evaluate this idea, we investigate the relevance of individual terms in the lexicon to each of the subcommunities. Table 3 shows the highest-ranking lexicon terms for each sub-community. These words align with the known focuses for the different groups; for example, the Pick-up Artists use terms such as IOI (indicators of interest) and SMV (sexual market value) and LTR (long-term relationship). To enable future reseach on these communities, we will publish the score of each term in the lexicon for each sub-community along with the lexicon.
The Spearman correlations of the word relevance across different subcommunities are shown in Figure 3. The Incels have negative correlations with all other communities (with a minimum $\rho=-0.240$ compared to the Men’s Rights Activists), showing that they are distinctive in their in-group language.
The largest correlation ($\rho=0.729$) is observed between the Pick-up Artists (PuA) and The Red Pill (TRP). This has been observed in prior work in sociology: Bachaud (2024) states that PuA terminology is plentiful in the TRP community, as they have a shared focus on seduction. The MGTOWs have a fairly strong correlation with TRP ($\rho=0.615$) and MRA ($\rho=0.654$). This aligns with their historical connections: Bachaud (2024) states that MGTOW was created by a group of MRAs and that their ideology is “close to both Red Pillers (critical about female nature) and MRAs (denouncing feminism and structural biases against men)”. A much weaker correlation is observed between MGTOW and PuA ($\rho=0.282$), which can be explained by their conflicting ideologies: PuAs are intent on pursuing women, whereas the main premise of MGTOW is to eschew contact with women. This captures an interesting insight into the relationship between PuAs, MGTOWs and TRPs: non-overlapping parts of the TRP interests are shared by both other groups.
The Criticism groups, who are reformed manosphere members, also correlate positively with MGTOW, MRA, TRP and PuA. Investigating their most relevant words (Table 3), we note that it mostly lists references to the manosphere itself.
Figure 3: Spearman correlations between in-group language of different manosphere sub-groups.
Group Top words
MGTOW thot, awalt, gyow, thots, simps, simp, beta, mgtow, cucked, alpha, cuck, bitches
Incels normie, blackpill, jfl, volcel, suifuel, braincels, normies, foids, incels, foid, blackpilled, stacy
Mental health hurt, failure, personality, killing, cope, harm, kill, hit, tbh, sa, hurt, dumb, beat
TRP trp, smv, rp, ltr, oneitis, alpha, beta, awalt, redpill, bp, asktrp, hamstering
MRA rape, raped, assault, patriarchy, female, abused, mras, dumb, hurt, mgtow, raping, fgm
PuA ioi, trp, smv, ltr, oneitis, alpha, iois, personality, beta, seddit, daygame, gf
Criticism trp, incels, manosphere, alpha, beta, rp, dumb, personality, asshole, redpill, redpilled, mgtow
Table 3: Highest relevance lexicon terms per category. Words in bold are novel contributions, i.e. not included in the lexicons in Table 1.
7 Related work
A number of interesting prior works relate to ours. We highlight some of their findings here.
Chancellor et al. (2016) investigate linguistic variation and the processes by which existing words are modified to create neologisms in the context of pro-anorexia groups, finding that lexical innovations are associated with increased engagement and more extreme content in the context of pro-anorexia online communities. Danescu-Niculescu-Mizil et al. (2013) explored the role of linguistic evolution in groups, using snapshot n-gram language models to measure the deviation of a user from the community. They illustrate that the non-adoption of lexical innovations is predictive of user churn in online communities. Zhang et al. (2017) characterises how distinctive and temporally dynamic online communities are in their language, finding that communities with more dynamic and distinctive language are likely to have higher retention of members.
Finally, Stewart and Eisenstein (2018) and Del Tredici and Fernández (2018) investigate factors that influence the dissemination of new lexical innovations, finding that linguistic and social factors both play a role in this process.
8 Conclusion
In this paper, we introduced and evaluated a novel framework for lexicon induction based on pretrained word embeddings. We compare several approaches based on two different basic models to methods from prior work, finding that the proposed methods achieve better scores on a newly crafted, expert-validated test set. A set of 455 new manosphere terms are to be released based on this analysis, with relevance scores for each manosphere sub-group.
Our work further provides novel insights into in-group language, including its embedding stability and correlation to violent language. The investigation into sub-group linguistic specialisation provides insight into the relationships between the manosphere subcommunities, which is supported by claims from prior work in sociology.
To our knowledge, this is the first work to incorporate social and temporal information to induce lexicons of in-group language. As the methodology is not specific to a particular group or language, we hope that it will be applied to create lexical resources for other communities in future.
Ethical considerations
The experiments in this paper make use of data developed in prior work. Ribeiro et al. (2021) state that the ethical standard guidelines of Rivers and Lewis (2014) were followed in the construction of this dataset, which includes not making any attempt to de-anonymize or link users across platforms. Even so, we recognise that the users did not consent to having their data analysed for the purpose of scientific experiments. However, given the real and current threat posed by this group, we view it is an acceptable use of the data. To limit the exposure of individuals, we do not provide any quotations or usernames and all analyses are performed at an aggregate level.
As we are mindful of potential mental health repercussions of this work to collaborators, participants (e.g. annotators) who were exposed to texts originating from manosphere communities were encouraged to attend a support group for researchers who work with extreme content.
Limitations
There are a number of simplifying assumptions that are made in this work. All models evaluated (LISTN and baselines) provide only one score per word, and hence do not account for sense variation or polysemy. We further include only single-word lexical variations in our evaluation and exclude non-alphanumeric characters. Our motivation for these choices is provided in Section 4.1.
The LISTN and word2vec models are trained on data from 2018. Given the dynamic nature of in-group language, this limits the applicability of the lexicon. Given recent limitations placed on Reddit data access, the Ribeiro et al. (2021) dataset still constitutes one of the largest and most varied existing manosphere datasets. Future work may look at applying these methods to more modern datasets (possibly in the context of specialised forums like incels.is); however, the evaluations in this work serve to validate our lexicon induction approach.
Appendix A Preliminary evaluation
Results for the first step of the evaluation are shown in Table 4.
AP AUROC
Baselines
PPMI 0.2162 0.3576
PPMI-subreddit 0.2564 0.4293
PPMI-category 0.2103 0.3470
NPMI 0.2864 0.4171
NPMI-subredit 0.3377 0.5322
NPMI-category 0.2748 0.4114
LISTN-CA
Community centroid 0.6147 0.7344
Category centroid 0.5577 0.6997
Subreddit centroid 0.3987 0.6287
Cluster-5 0.4302 0.637
Cluster-100 0.3399 0.5554
Bootstrap 0.2811 0.4960
Bias 0.5690 0.7065
LISTN-C
Community centroid 0.6147 0.7288
Category centroid 0.6809 0.7272
Subreddit centroid 0.5461 0.5891
Cluster-5 0.491 0.6827
Cluster-100 0.4007 0.6184
Bootstrap 0.5276 0.5077
Bias 0.6015 0.5382
Table 4: Results for preliminary experiments, used to construct the gold label test set.